OpenStudy (shelby1290):

State the missing factor 4n^2+8n-60=(?)(n+5)

OpenStudy (anonymous):

what times \(n\) is \(4n^2\)?

OpenStudy (shelby1290):

4n?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and what times \(5\) gives \(-60\)?

OpenStudy (shelby1290):

-12

OpenStudy (anonymous):

ok put them together are you are done

OpenStudy (shelby1290):

(4n-12)

OpenStudy (anonymous):

yup

OpenStudy (shelby1290):

at the back of my textbook the answer is 4(n-3)

OpenStudy (anonymous):

not really factored completely that would be \[4(n-3)(n+5)\]

OpenStudy (anonymous):

yeah because the terms of \(4n-12\) have a common factor of 4

OpenStudy (shelby1290):

so you divide (4n-12) by 4 to get 4(n-3) ?

OpenStudy (anonymous):

that is not "division" that is factoring, aka the distributive property \[4\times x-4\times 3=4(x-3)\]

OpenStudy (shelby1290):

ohh okay